# Distributed Online Feedback Optimization for Real-time Distribution System Voltage Regulation

### v2

@@ II-D Proposed distributed approach
The two-metric approach in Section II-C does not necessarily lead to algorithm convergence since it is not in general a descent iteration [13]. To ensure convergence, the projection in (3) should use the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ due to the scaling matrix $\mathbf{X}^{-1}$ [13], that is:
@@
$\mathbf{q}^{k+1}\leftarrow\underset{\mathbf{u}\in\mathcal{X}^{k}}{\operatorname{argmin}}\quad\frac{1}{2}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})^{\intercal}\mathbf{X}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})$ (4)
@@
This however contradicts the distributed communication architecture of the overall OFO algorithm since, unlike the Euclidean projection (equivalent to the case where $\mathbf{X}$ is the identity matrix), solving (4) again requires global information. As a solution, [12] proposed dualization of the DER capacity limits, combined with the Euclidean projection, to ensure the DER constraint satisfaction. However, as demonstrated both in their numerical result and our simulation, this approach requires an excessively large number of iterations to converge, significantly hindering its performance in online implementation. To tackle this challenge, this paper proposes an iterative algorithm to solve (4) that requires only local information and is implemented in an online feedback-based manner. This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (4) to generate feasible setpoints. Our key observation is that given a feasible $\mathbf{u}^{\tau}\in\mathcal{X}^{k}$, the gradient of the objective function in (4) at $\mathbf{u}^{\tau}$ is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm. The proposed algorithm is based on the non-scaled Euclidean gradient projection and proceeds as:
@@
    Implement $\mathbf{q}^{k}_{\epsilon}:=\mathbf{q}^{k}+\epsilon(\mathbf{\dot{q}}^{k+1}-\mathbf{q}^{k})$ and measure the voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$. Estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ as $\mathbf{\tilde{v}}^{k}$ + ($\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$-$\mathbf{\tilde{v}}^{k}$)/$\epsilon$.
@@
    $\mathbf{u}^{\tau+1}\leftarrow[\mathbf{u}^{\tau}-\alpha^{u}(\mathbf{\tilde{v}}(\mathbf{u}^{\tau})-\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1}))]_{\mathbf{\underline{q}}^{k}}^{\mathbf{\overline{q}}^{k}}$ (5)
@@
In the first step, $\epsilon$ represents a small exploration parameter, which is similar to that in model-free feedback optimization [14]. Since $\mathbf{\dot{q}}^{k+1}$ can be pretty far away from the feasible set, it might be impossible to implement it. Therefore, we implement $\mathbf{q}^{k}_{\epsilon}$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ based on it. This could still result in minor temporary DER capacity limit violations, which should be manageable by the DERs. To ensure strict capacity limit satisfaction, the original feasible set can be slightly deflated [14]. The initial feasible $\mathbf{u}^{\tau}$ in the second step can be chosen as $\mathbf{q}^{k}$. Figure 2 shows the implemented reactive power setpoints and communication requirements in an outer iteration.
Figure 2: Implemented reactive power setpoints and communication requirements in an outer iteration for the proposed strategy.
II-E Convergence analysis
The convergence analysis of the nested algorithm is based on the following two assumptions. First, we assume a linear relation holds between voltages and reactive power injections. That is, $\mathbf{v}=\mathbf{v}^{0}+\mathbf{X}\mathbf{q}$, where the constant term $\mathbf{v}^{0}$ captures the impact of other disturbances. Based on the first assumption, $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ is the exact gradient of the objective function in (4) at $\mathbf{u}^{\tau}$. The inner loop then represents a standard Euclidean gradient projection algorithm to solve the quadratic program (4). Consequently, $\mathbf{u}^{\tau}$} is a converging sequence by invoking Proposition 2.3.2 in [13] and $\alpha^{u}$ chosen with the rule therein. Second, we assume the inner iterations indeed converge after $T$ iterations. The overall nested algorithm thus represents a scaled PDGP algorithm to solve (1). By a transformation of variables defined by: $\mathbf{q}^{\prime}=\mathbf{X}^{\frac{1}{2}}\mathbf{q}$ and $\mathbf{v}^{\prime}=\gamma\mathbf{v}$ where $\gamma=\frac{\alpha^{d}}{\alpha}$ is the ratio between the dual and primal step sizes, the problem (1) can be cast as a quadratic program in the space of $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$. Likewise, the overall nested algorithm can be cast as an equivalent unscaled PDGP algorithm in $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$ [13]. Invoking Theorem 3 in [5] and with $\alpha$ chosen sufficiently small within the upper bound defined therein, the overall algorithm then converges to the optimizer of the regularized saddle-point problem of (1) at a Q-linear rate.
@@ II-E Convergence analysis
Remark: First, while the linear voltage-reactive power relation is assumed to synthesize the controller and analyze its convergence, AC power flow is run in the simulation to calculate the real voltages. Such an assumption was also made in the proofs of [1, 11]. Second, a large $T$ is not necessary in the simulation. We use $T=4$ so the inner loop might not converge. However, the algorithm performs well in the simulation. Notably, a similar assumption was made in [1].
III Simulation results
@@
In the numerical simulation, we study a 100% penetration scenario of photovoltaics (PVs) in a 96-node system, which experiences overvoltage issues. Details of the system and the generation and load data with a 6-second resolution can be found in [7] and references therein. The PV inverters are assumed to be oversized by 20% to provide sufficient reactive power. The reactive power cost matrix $\mathbf{C}$ is chosen as the identity matrix. The lower and upper voltage limits are 0.95 and 1.05 pu, respectively. The unit for power is $kW/kVar$. The step sizes are chosen with a trial-and-error strategy [12] as $\alpha^{d}=10^{6}$, $\alpha=2\times 10^{-4}$, and $\alpha^{u}=10^{2}$. The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 1%. The primal and dual regularization is used to establish the theoretical convergence result. The impact of some small regularization parameters on the numerical results is negligible. In the simulation, we set them as $10^{-4}$. For each generation and load data point, an OFO iteration is run which yields tentative reactive power setpoints, followed by one exploration step and 4 inner feedback iterations of (5), i.e. we assume that every setpoint in Fig. 2 is implemented for 1 second. Finally, the nested algorithm requires only basic arithmetic operations and is computationally very efficient. For a series implementation of a 4-hour simulation, i.e. in total 14400 iterations, it takes 7.6 seconds, averaging 0.5 millisecond per iteration.
@@
Figure 3: Voltage profiles over 1000 iterations for several buses in the static case using the centralized and the proposed approaches. We count both the outer and inner iterations for our proposed approach.
@@
Figure 4: Reactive power profiles after 1000 iterations for the centralized and the proposed approaches.
Figure 5: Voltage profiles for the most remote bus in the dynamic simulation. We define the average voltage violation (AVV) as $\operatorname{AVV}=\frac{1}{K}\sum_{k=1}^{K}[v^{k}-\overline{v}]^{+}$, where $k$ is the iteration index and $K$ is the number of iterations. The AVV values for the various approaches are: $1.5\times 10^{-2}$ (no control), $8.3\times 10^{-5}$ (centralized), $2.5\times 10^{-3}$ (two-metric), $2.1\times 10^{-4}$ (proposed), $3.3\times 10^{-3}$ (Tang 2024 [11]), $8.6\times 10^{-4}$ (Qu 2020 [12]).
@@
In the static case, the generation and load data are kept unchanged to study the convergence of the nested approach. Figure 3 shows that the proposed approach suppresses the initial voltage limit violations quickly and converges after around 300 total iterations, i.e. 50 outer OFO iterations. The approach also utilizes the available grid capacity as efficiently as the centralized one. Figure 4 compares the converged reactive power setpoints, which shows that the centralized and the proposed approaches converge to the same solution. This is reasonable since the gradient scaling only alters the converging trajectory, but does not change the solution to (1).

### v3

@@ III-B Proposed distributed approach
It has been shown in [25] that, to ensure convergence, the projection should use the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ due to the scaling matrix $\mathbf{X}^{-1}$ [25], that is:
@@
$\mathbf{q}^{k+1}=\underset{\mathbf{u}\in\mathcal{X}^{k}}{\operatorname{argmin}}\quad\frac{1}{2}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})^{\intercal}\mathbf{X}(\mathbf{u}-\mathbf{\dot{q}}^{k+1}).$ (13)
This however contradicts the distributed communication architecture of the overall OFO algorithm since, unlike the Euclidean projection, solving (13) again requires global information. As a solution, [24] proposed dualization of the DER capacity limits, combined with the Euclidean projection, to ensure the DER constraint satisfaction. However, this approach requires an excessively large number of iterations to reach convergence, significantly hindering its performance in online implementation.
Figure 4: Illustration of the approximation.
@@
To tackle this challenge, this paper proposes an iterative algorithm to solve (13) that requires only local information and is implemented in an online feedback-based manner. This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (13) to generate feasible setpoints. Our key observation is that given a feasible $\mathbf{u}^{\tau}\in\mathcal{X}^{k}$, the gradient of the objective function in (13) at $\mathbf{u}^{\tau}$ is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ according to (2) assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm, i.e.
$\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})\approx\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1}).$ (14)
The proposed algorithm is based on the non-scaled Euclidean gradient projection and proceeds as:
@@
    Implement
    $\mathbf{q}^{k}_{\epsilon}=\mathbf{q}^{k}+\epsilon(\mathbf{\dot{q}}^{k+1}-\mathbf{q}^{k}),$ (15)
    and measure the voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$. Estimate
    $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})\approx\mathbf{\tilde{v}}^{k}+(\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})-\mathbf{\tilde{v}}^{k})/\epsilon.$ (16)
@@
    $\mathbf{u}^{\tau+1}=[\mathbf{u}^{\tau}-\alpha^{u}(\mathbf{\tilde{v}}(\mathbf{u}^{\tau})-\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1}))]_{\mathbf{\underline{q}}^{k}}^{\mathbf{\overline{q}}^{k}}.$ (17)
Remark 1.
@@
In the first step, $\epsilon$ represents a small exploration parameter, which is similar to that in model-free feedback optimization [29, 30, 31, 32]. Since $\mathbf{\dot{q}}^{k+1}$ can be pretty far away from the feasible set, it might be impossible to implement it. Therefore, we implement $\mathbf{q}^{k}_{\epsilon}$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ based on it. This is illustrated in Fig. 4. This could still result in minor temporary DER capacity limit violations, which should be manageable by the DERs. To ensure strict capacity limit satisfaction, the original feasible set can be slightly deflated [32]. The initial feasible $\mathbf{u}^{\tau}$ in the second step can be chosen as $\mathbf{q}^{k}$. The online implementation of the overall nested algorithm is summarized in Algorithm 1.
Data: $\mathbf{q}^{0}$
Result: $\mathbf{q}^{k}$, $k\geq 0$; $\mathbf{u}^{\tau}$, $\tau\leq T$
1 Set k = 0;
2 while $k\geq 0$ do
3       Implement $\mathbf{q}^{k}$ and collect voltage measurement $\tilde{\mathbf{v}}(\mathbf{q}^{k})$;
4       Communicate $\mathbf{q}^{k}$ between physical neighbours and calculate $\dot{\mathbf{q}}^{k+1}$ via (11);
5       Implement $\mathbf{q}^{k}_{\epsilon}$, measure voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$;
6       Set $\tau=0$;
7       while $\tau<T$ do
8             Implement $\mathbf{u}^{\tau}$ and measure voltage $\tilde{\mathbf{v}}(\mathbf{u}^{\tau})$ (choose $\mathbf{u}^{0}$ as $\mathbf{q}^{k}$);
9             Calculate next implementable $\mathbf{u}^{\tau+1}$ via (17);
10             Set $\tau=\tau+1$;
11
12       end while
13      Set $\mathbf{q}^{k+1}=\mathbf{u}^{T}$;
14       Set $k=k+1$.
15 end while
Algorithm 1 Nested feedback optimization algorithm
III-C Convergence analysis
The convergence analysis of the nested algorithm is based on the following two assumptions.
Assumption 1.
The linear relation (2) holds between voltages and reactive power injections.
Based on the first assumption, $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ is the exact gradient of the objective function in (13) at $\mathbf{u}^{\tau}$. The inner loop then represents a standard Euclidean gradient projection algorithm to solve the quadratic program (13). Consequently, $\mathbf{u}^{\tau}$} is a converging sequence by invoking Proposition 2.3.2 in [25] with $0<\alpha^{u}<2/\lambda_{max}(\mathbf{X})$ where $\lambda_{max}(\mathbf{X})$ is the largest eigenvalue of $\mathbf{X}$.
Assumption 2.
The inner iterations reach convergence after $T$ iterations.
The overall nested algorithm thus represents a scaled PDGP algorithm to solve (6). By a transformation of variables defined by: $\mathbf{q}^{\prime}=\mathbf{X}^{\frac{1}{2}}\mathbf{q}$ and $\mathbf{v}^{\prime}=\gamma\mathbf{v}$ where $\gamma=\sqrt{\frac{\alpha^{d}}{\alpha}}$ is the squared root of the ratio between the dual and primal step sizes, the problem (6) can be cast as an equivalent quadratic program in the space of $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$ as in (18).
$\underset{\mathbf{q^{\prime}}}{\operatorname{minimize}}$ $\quad\frac{1}{2}\mathbf{q^{\prime}}^{\intercal}\mathbf{X}^{-\frac{1}{2}}\mathbf{C}\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}},$ (18a)
$\operatorname{s.t.}$ $\quad\gamma\mathbf{\underline{v}}\leq\mathbf{v^{\prime}}\leq\gamma\mathbf{\overline{v}}:\boldsymbol{\mu^{\prime}},\boldsymbol{\lambda^{\prime}},$ (18b)
$\mathbf{\underline{q}}\leq\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}}\leq\mathbf{\overline{q}}.$ (18c)
The unscaled regularized PDGP algorithm to solve (18) includes the following iterations:
$\boldsymbol{\lambda^{\prime}}^{k+1}=[\boldsymbol{\lambda^{\prime}}^{k}+\alpha(\gamma\mathbf{\tilde{v}}-\gamma\mathbf{\overline{v}}-r^{d}\boldsymbol{\lambda^{\prime}}^{k})]^{+},$ (19a)
$\boldsymbol{\mu^{\prime}}^{k+1}=[\boldsymbol{\mu^{\prime}}^{k}+\alpha(\gamma\mathbf{\underline{v}}-\gamma\mathbf{\tilde{v}}-r^{d}\boldsymbol{\mu^{\prime}}^{k})]^{+},$ (19b)
$\mathbf{q^{\prime}}^{k+1}=\underset{\mathbf{q^{\prime},\mathbf{\underline{q}}\leq\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}}\leq\mathbf{\overline{q}}}}{\operatorname{argmin}}\lVert\mathbf{q^{\prime}}-\big{[}\mathbf{q^{\prime}}^{k}-\alpha[\mathbf{X}^{-\frac{1}{2}}$ (19c)
$\qquad\cdot\mathbf{C}\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}}^{k}+\gamma\mathbf{X}^{\frac{1}{2}}(\boldsymbol{\lambda^{\prime}}^{k+1}-\boldsymbol{\mu^{\prime}}^{k+1})+r^{p}\mathbf{q^{\prime}}^{k}]\big{]}\rVert_{2}^{2}.$
which are equivalent to (9a)-(9b) and (11)-(13) [25, Eqs. (2.37) and (2.38)]. Invoking Theorem 3 in [2] with $\alpha$ chosen using [2, Eq. (34)], (19) then converges to the optimizer of the regularized saddle-point problem of (18). Consequently, our proposed nested algorithm also converges to the optimizer of the regularized saddle-point problem of (6).
Remark 2.
@@ III-C Convergence analysis
First, while the linear voltage-reactive power relation is assumed to synthesize the controller and analyze its convergence, AC power flow is run in the simulation to calculate the actual voltages. Such an assumption was also made in the proofs of [3, 23]. Second, a large $T$ is not necessary in the simulation. We use $T=4$ so the inner loop might not converge. However, the algorithm performs well in the simulation. Notably, a similar assumption was made in [3].
IV Case study
IV-A Case description
Figure 5: A 96-bus test system from Simbench [33].
In the numerical simulation, we study a 100% penetration scenario of photovoltaics (PVs) in a 96-bus system, i.e. we assume every node has a PV installation. The test system is adopted from Simbench [33] and is shown in Fig. 5, which is expected to experience overvoltage issues. The network topology and impedance data are the same as those in the original dataset. For each bus, a PV installation with a randomly generated DC capacity from 3-10 kW is assumed. The PV inverters are assumed to be oversized by 20% to provide sufficient reactive power. For the dynamic case study, generation and load data with a 6-second resolution are used. The PV profiles are linearly interpolated using the HelioClim-3 dataset [34] with a 1-minute resolution. Base load profiles are obtained by aggregating 1-second resolution data from the ECO dataset [35]. AC power flow problems are solved with a high-performance library PowerGridModel [36].
@@ IV-A Case description
The reactive power cost matrix $\mathbf{C}$ is chosen as the identity matrix. The lower and upper voltage limits are 0.95 and 1.05 pu, respectively. The unit for power is kW/kVar. The step sizes are chosen with a trial-and-error strategy [24] as $\alpha^{d}=10^{6}$, $\alpha=5\times 10^{-4}$, and $\alpha^{u}=10^{2}$. The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 0.6%. The primal and dual regularization is used to establish the theoretical convergence result. The impact of some small regularization parameters on numerical results is negligible. In the simulation, we set them as $10^{-4}$. For each generation and load data point, an OFO iteration is run which yields tentative reactive power setpoints, followed by one exploration step and 4 inner feedback iterations of (17), i.e. we assume that every setpoint in Algorithm 1 is implemented for 1 second. Finally, the nested algorithm requires only basic arithmetic operations and is computationally very efficient. For a series implementation of a 4-hour simulation, i.e. in total 14400 iterations, it takes 8.1 seconds, averaging 0.56 milliseconds per iteration.
@@
Figure 6: Voltage profiles over 500 iterations for several buses in the static case using the centralized and the proposed approaches. We count both the outer and inner iterations for our proposed approach.
@@
Figure 7: Reactive power profiles after 500 iterations for the centralized and the proposed approaches.
(a) No control
(b) Centralized
(c) Proposed
(d) Two metric
(e) Tang 2024 [23]
(f) Qu 2020 [24]
(g) Droop control [37, 38]
(h) Summary for bus 95
Figure 8: Voltage profiles under different approaches in the dynamic simulation.
(a) Centralized
(b) Proposed
(c) Two metric
(d) Tang 2024 [23]
(e) Qu 2020 [24]
(f) Droop control [37, 38]
Figure 9: Reactive power profiles under different approaches in the dynamic simulation.
IV-B Static case results
@@
In the static case, the generation and load data are kept unchanged to study the convergence of the nested approach. The simulation uses data at 12:00. Figure 6 shows that the proposed approach suppresses the initial voltage limit violations quickly and converges after around 200 total iterations, i.e. 33 outer OFO iterations. The approach also utilizes the available voltage limit as efficiently as the centralized one. Figure 7 compares the converged reactive power setpoints, which shows that the centralized and the proposed approaches converge to the same solution. This is reasonable since the gradient scaling only alters the converging trajectory, but does not change the solution to (6).
IV-C Dynamic case results
In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online to track the time-varying optimizers of (6) without waiting for them to converge. Figure 8a shows significant voltage limit violations without control. Figures 8b and 8c demonstrate that the centralized and our proposed distributed nested feedback optimization approaches both successfully enforce the voltage limit, although temporary voltage limit violations are inevitable due to the corrective nature of these algorithms. The simple two-metric approach and the heuristic approach in [23] result in prolonged voltage limit violations as seen in Figs. 8d and 8e, respectively. Figure 8f shows that the distributed approach in [24] does not track the time-varying optimizers well and experiences oscillations at the end of the simulation period. The discrepancy between the actual setpoints and algorithm iterates (without projection) potentially renders the algorithm slow and unstable. Figure 8g shows that the local droop control approach does not enforce the voltage limit due to lack of coordination. Finally, for clarity, Fig. 8h presents voltages for the most remote, and therefore, the most sensitive bus (bus 95) under different approaches.
Furthermore, Fig. 9 shows reactive power profiles under different approaches. Shown in Figs. 9a and 9b, our proposed distributed approach achieves similar profiles with the centralized one, with an average setpoint deviation of $-1.3\times 10^{-3}$ kVar. The two-metric approach and the heuristic approach do not work well in this case study. The approach in [24] experiences oscillations. The local droop control approach does not in theory or practice track the optimizers of (6).
Finally, we define the average voltage violation (AVV) in (20), where $k$ is the iteration index and $K$ is the total number of iterations.
$\operatorname{AVV}=\frac{1}{K}\sum_{k=1}^{K}\left([\tilde{v}^{k}-\overline{v}]^{+}+[\underline{v}-\tilde{v}^{k}]^{+}\right).$ (20)
TABLE I: Average voltage violations for various approaches.
Approach AVV (pu) w.r.t. centralized
No control $1.5\times 10^{-2}$ 192x
Centralized $8.0\times 10^{-5}$ 1x
Proposed $1.6\times 10^{-4}$ 2x
Two-metric $2.5\times 10^{-3}$ 31x
Tang 2024 [23] $3.3\times 10^{-3}$ 41x
Qu 2020 [24] $8.6\times 10^{-4}$ 11x
Droop control [37, 38] $4.5\times 10^{-4}$ 6x